### v2

@@ 5 Evaluation
The framework is evaluated through a design science experiment involving prospective users to assess its practical usability and real-world applicability. To contextualize this evaluation, we consider the ORKG as a representative environment for scientific content creation. Following the goal definition template by Wohlin et al. [42], we define the experiment’s objective as follows:
@@
Research goal: Analyze the developed HITL tool with the purpose of supporting systematic content creation for scientific KGs in terms of usability (effectiveness, efficiency, satisfaction) from the perspective of potential future users familiar with the ORKG in the context of the Open Research Knowledge Graph. The goal leads to the research question: {mdframed} Research question: How does the developed HITL tool impact the usability of systematic content creation for scientific knowledge graphs, from the perspective of potential future users familiar with the ORKG? The experiment aims to assess the practical usability of the developed HITL tool for creating scientific content for knowledge graphs. To this end, we collaborated with ORKG curators who have previously created comparisons as part of literature reviews. Each participant will first subjectively assess the time they spent on manually creating one of their existing comparisons, reflecting state-of-the-art research in a specific domain. They will then recreate the same comparison using the HITL tool, without reviewing the original version to avoid bias.
The developed framework is the independent variable, while effectiveness, efficiency, and satisfaction are the dependent variables for usability. Effectiveness is measured by participants’ perception and objectively by the number of identified papers, along with the accuracy of extracted data compared to the established comparison. Efficiency is assessed through participants’ perception and objectively by the time taken to create the comparison. Satisfaction is evaluated subjectively after each stage and overall using the System Usability Scale (SUS) [24] at the experiment’s conclusion.
With particpants’ consent, both screen and audio were recorded during the sessions to enable detailed post-analysis. The experiment is described in a document [16] provided to participants at the start, outlining the study context and presenting title and description of their most recent comparison. The HITL tool runs locally on a computer provided by the experimenter. Sessions are conducted either in person or remotely via TeamViewer, ensuring a consistent experience. A questionnaire collects demographic data, ORKG usage frequency, and consent. Participants rate their agreement with statements regarding the tools’ usefulness and user experience using a Likert scale. The SUS is administered at the end of each session, alongside open-ended questions for additional qualitative feedback.
@@
    Participants proceed through the two stages of the workflow, responding to related questionnaire items while engaging in the Think-Aloud method [34].
@@
The raw data [16] is published on Zenodo. Nine participants, including external curators and internal members, took part in the experiment, primarily from Semantic Web technologies or neuro-symbolic AI fields. Participants reported varied ORKG usage frequencies, ranging from weekly to monthly, with experience spanning from a few months to several years, reflecting different levels of expertise with the system. Figure 6 presents the participants’ responses regarding subjective efficiency and effectiveness, showing strong agreement that the tool makes corpus creation and knowledge extraction ‘faster’ and ‘easier’, and a weaker agreement to its ability to render ‘better results’.
@@
Figure 7 represents the resulting SUS scores. We achieve an overall SUS score of 84.17 ($min=75$, $max=90$, $\sigma=4.86$), regarded as A+ score – the highest usability rating [24]. Participants expressed general satisfaction with the tool, with comments like “The workflow should be integrated to ORKG as soon as possible. I will like to use it” and “Very helpful system!!!”. Specific feedback highlighted features like the “Keywords extraction option is an amazing idea”. Suggestions for improvement included adding “Properties should have descriptions (for more accurate extraction)” and “focus could be on extracting entities that align with the similar papers already available in the ORKG”.
@@
The reported durations for the participants’ original comparison varied significantly, with the fastest comparison taking approximately 4 hours to complete, while the longest required around 2 weeks. Completing the workflow with our framework – from defining a research interest based on their original comparison to successfully creating a new ORKG comparison – took an average of 1,480 seconds (24:40 minutes), with a minimum of 1,023 seconds (17:03 minutes) and a maximum of 1,996 seconds (33:16 minutes), as detailed in Figure 8.
@@
QIP.
The QIP methodology showed that curators are a valuable user group for guiding tool development. To improve CSV-based integration of literature tables, enhanced semantification is currently being developed to better align entities within the ORKG [17].
@@
Based on Wohlin et al. [42], we identify potential threats to the validity of our experiment. Conclusion Validity. Participants may subconsciously replicate elements from their original ORKG comparison. To reduce this risk, we did not inform them that the task was based on their previous work. Participants came from diverse research fields but worked on their own prior content, minimizing the impact of disciplinary differences while supporting broader generalizability. Internal Validity. Participants may have been predisposed to emphasize the tool’s benefits in comparison to doing it without the tool, particularly ease and speed. As part of the participants were ORKG curators, their financial affiliation could have introduced a bias toward favorable feedback, given the tool’s relevance to their ongoing work. Construct Validity. To capture genuine reactions and difficulties, we applied the Think-Aloud Method [34], encouraging participants to verbalize thoughts during each task. This helped ensure feedback reflected actual experiences with the tool. External Validity. With only nine participants, generalizability is limited. Nonetheless, the range of disciplinary backgrounds offers insight into the tool’s broader applicability. Future evaluations should involve more diverse and numerous participants to strengthen external relevance.
@@ 6 Discussion and Future Work
We interpret the mean SUS score using Bangor et al.’s [6] adjective scale, where the tool’s score of 84.17 ranks at the upper boundary of “Good”, and an A+ in the Lewis and Sauro’s benchmarks [24], indicating a positive user experience. Participants described the tool as intuitive, easy to navigate, and effective for task execution. The tool significantly reduced the time needed to create content in the ORKG compared to participants’ prior methods. While this shows strong potential for streamlining workflows, it does not automatically ensure higher-quality outputs. Participants appreciated the simplified corpus creation and extraction process but expressed concerns about accuracy and over-reliance on automated suggestions. Those who invested more time refining the corpus reported better outcomes, suggesting deeper engagement improves quality. Still, not all agreed the tool consistently produced better results, exposing a gap between automation and human expectations. This disparity points to a key challenge in HITL systems: balancing automation with human judgment to ensure both efficiency and high-quality, contextually relevant results. Addressing this requires enhancing the extraction process and incorporating iterative feedback loops to refine accuracy. Although the experiment aimed to replicate a prior comparison, measuring replication quality through the number of identified papers proved unfeasible as an indicator of replication quality, underscoring the challenge of replicating structured research output in a HITL system. Unlike automated systems, our approach relies on user insights and interpretations, making the process subjective and context-dependent.
Requirements.
@@
The tool satisfies most core requirements defined for the HITL framework. It guides users from defining a research interest to extracting and editing data, with CSV export for seamless integration into scholarly KGs like the ORKG. The modular pipeline supports extensibility, adaptability, and long-term semantic preservation. However, key limitations remain. Entity linking (FR12) is not yet implemented, leaving extracted entities disconnected from existing KG resources. Incorporating entity alignment and exporting in formats like JSON-LD or RDF would improve semantic interoperability. Support for processing existing corpora (FR7) is also insufficient. Currently, users must manually import PDFs and resolve potential errors independently. Integration with tools like Zotero could streamline corpus handling and reduce user friction. Additionally, performance degrades with large datasets (NR4), as processing occurs sequentially per document. Future work should focus on performance optimization and parallelization strategies. Participants also requested more control over the data model (FR8). Allowing users to define domain-specific properties and extraction instructions would improve precision and adaptability across fields. This aligns with feedback on the need for more customizable and explainable outputs. Crucially, while automation accelerates literature processing, LLM-based extractions still require human validation (FR10, FR11). Without critical oversight, there’s a risk of misinterpreting or blindly trusting generated results. This underscores the ethical implications of AI-assisted research, especially risking the loss of genuine understanding and knowledge internalization.
Future Work.
@@
Future development should allow storing and refining outputs with customizable data models, enabling users to iteratively improve extraction accuracy. Users can selectively re-run extractions on specific cells while preserving high-quality results, enhancing the HITL process. Semantic structuring, like linking content to knowledge graph entities and supporting interoperable formats, will strengthen the system’s utility. Ongoing improvements in reference manager integration, scalability, and model configurability will stabilize and enrich the tool. Lastly, research should focus on user guidance for validating and interpreting extractions, e.g., by highlighting source passages in PDFs, ensuring human oversight in AI-driven workflows.
@@ 7 Conclusion
Our work presents a semi-automated solution for structuring scientific content specifically for integration into the ORKG. Aligned with the quality improvement (QIP) paradigm, we designed a HITL workflow consisting of two stages, Creating a scientific corpus of literature and Knowledge extraction and validation. Our modular framework implemented this workflow in a Jupyter notebook environment and assists in identifying relevant publications, extracting key information, and structuring knowledge into reusable formats for integration into scientific knowledge graphs, exemplified by the ORKG.
By automating repetitive tasks and allowing researchers to interactively refine outputs, the tool reduces the manual effort typically involved in processing large volumes of literature, moving towards making large-scale literature reviews more efficient and enabling researchers to keep pace with the expanding body of scientific knowledge. While the evaluation results showed the tool is highly usable and accelerates the process, we identified areas for improvement. Additional support is needed to engage users in the knowledge extraction process, alongside additional feedback loops to enhance the quality of extracted information.

### v3

@@ 5 Evaluation
The framework ExtracTable is evaluated through a design science experiment involving prospective users to assess its practical usability and real-world applicability. To contextualize this evaluation, we consider the ORKG as a representative environment for scientific content creation. Following the goal definition template by Wohlin et al. [42], we define the experiment’s objective as follows:
@@
Research goal: Analyze ExtracTable with the purpose of supporting systematic content creation for scientific KGs in terms of usability (effectiveness, efficiency, satisfaction) from the perspective of potential future users familiar with the ORKG in the context of the ORKG. The goal leads to the research question: {mdframed} Research question: How does the developed tool ExtracTable impact the usability of systematic content creation for scientific knowledge graphs, from the perspective of potential future users familiar with the ORKG? The experiment aims to assess the practical usability of the developed tool ExtracTable for creating scientific content for knowledge graphs. To this end, we collaborated with ORKG curators who have previously created comparisons as part of literature reviews. Each participant will first subjectively assess the time they spent on manually creating one of their existing comparisons, reflecting state-of-the-art research in a specific domain. They will then recreate the same comparison using ExtracTable, without reviewing the original version to avoid bias. We used a within-subject design to assess usability and enable meaningful self-comparison across diverse domain experts.
The developed framework ExtracTable is the independent variable, while effectiveness, efficiency, and satisfaction are the dependent variables for usability. Effectiveness is measured by participants’ perception and objectively by the number of identified papers, along with the accuracy of extracted data compared to the established comparison. Efficiency is assessed through participants’ perception and objectively by the time taken to create the comparison. Satisfaction is evaluated subjectively after each stage and overall using the System Usability Scale (SUS) [25] at the experiment’s conclusion.
With particpants’ consent, both screen and audio were recorded during the sessions to enable detailed post-analysis. The experiment is described in a document [16] provided to participants at the start, outlining the study context and presenting title and description of their most recent comparison. ExtracTable runs locally on a computer provided by the experimenter. Sessions are conducted either in person or remotely via TeamViewer, ensuring a consistent experience. A questionnaire collects demographic data, ORKG usage frequency, and consent. Participants rate their agreement with statements regarding the tools’ usefulness and user experience using a Likert scale. The SUS is administered at the end of each session, alongside open-ended questions for additional qualitative feedback.
@@
    Participants proceed through the two stages of ExtracTable, responding to related questionnaire items.
@@
The raw data [16] is published on Zenodo. Nine participants took part, including external curators and internal ORKG team members not involved in the tool’s development, mostly from Semantic Web or neuro-symbolic AI fields. Participants reported varied ORKG usage frequencies, ranging from weekly to monthly, with experience spanning from a few months to several years, reflecting different levels of expertise with the system. Figure 6 presents the participants’ responses regarding subjective efficiency and effectiveness, showing strong agreement that the tool makes corpus creation and knowledge extraction ‘faster’ and ‘easier’, and a weaker agreement to its ability to render ‘better results’.
@@
Figure 7 represents the resulting SUS scores. We achieve an overall SUS score of 84.17 ($min=75$, $max=90$, $\sigma=4.86$), regarded as A+ score – the highest usability rating [25]. Participants expressed general satisfaction with ExtracTable, with comments like “The workflow should be integrated to ORKG as soon as possible. I will like to use it” and “Very helpful system!!!”. Specific feedback highlighted features like the “Keywords extraction option is an amazing idea”. Suggestions for improvement included adding “Properties should have descriptions (for more accurate extraction)” and “focus could be on extracting entities that align with the similar papers already available in the ORKG”.
@@
The reported durations for the participants’ original comparison varied significantly, with the fastest comparison taking approximately 4 hours to complete, while the longest required around 2 weeks. We acknowledge that these self-reported durations are prone to imprecision but still provide an indicative measure of the time involved. Completing the workflow with our framework – from defining a research interest based on their original comparison to successfully creating a new ORKG comparison – took an average of 1,480 seconds (24:40 minutes), with a minimum of 1,023 seconds (17:03 minutes) and a maximum of 1,996 seconds (33:16 minutes), as detailed in Figure 8.
@@
QIP. The QIP methodology showed that curators are a valuable user group for guiding tool development. To improve CSV-based integration of literature tables, enhanced semantification is currently being developed to better align entities within the ORKG [17].
@@
Based on Wohlin et al. [42], we identify potential threats to the validity of our experiment. Conclusion Validity. Participants may subconsciously replicate elements from their original ORKG comparison. To reduce this risk, we did not inform them that the task was based on their previous work. Participants came from diverse research fields but worked on their own prior content, minimizing the impact of disciplinary differences while supporting broader generalizability. Internal Validity. Participants may have been predisposed to emphasize the tool’s benefits in comparison to doing it without the tool, particularly ease and speed. As part of the participants were ORKG curators, their financial affiliation could have introduced a bias toward favorable feedback, given the tool’s relevance to their ongoing work. Additionally, self-reported durations for prior comparison creation, collected in unobserved settings, are prone to imprecision, adding uncertainty when comparing to the objectively timed tool-based workflow. Construct Validity. To capture genuine reactions and difficulties, we encouraged participants to comment on their experience during each task. This ensured feedback reflected their actual interactions with the tool. External Validity. With only nine participants, generalizability is limited. Nonetheless, the range of disciplinary backgrounds offers insight into the tool’s broader applicability. Future evaluations should involve more diverse and numerous participants to strengthen external relevance.
@@ 6 Discussion and Future Work
Experiment. We interpret the mean SUS score using Bangor et al.’s [6] adjective scale, where ExtracTable’s score of 84.17 ranks at the upper boundary of “Good”, and an A+ in the Lewis and Sauro’s benchmarks [25], indicating a positive user experience. Participants described ExtracTable as intuitive, easy to navigate, and effective for task execution. ExtracTable significantly reduced the time needed to create content in the ORKG compared to participants’ prior methods. While this shows strong potential for streamlining workflows, it does not automatically ensure higher-quality outputs. Participants appreciated the simplified corpus creation and extraction process but expressed concerns about accuracy and over-reliance on automated suggestions. Those who invested more time refining the corpus reported better outcomes, suggesting deeper engagement improves quality. Still, not all agreed the tool consistently produced better results, exposing a gap between automation and human expectations. This disparity points to a key challenge in HITL systems: balancing automation with human judgment to ensure both efficiency and high-quality, contextually relevant results. Addressing this requires enhancing the extraction process and incorporating iterative feedback loops to refine accuracy. Although the experiment aimed to replicate a prior comparison, measuring replication quality through the number of identified papers proved unfeasible as an indicator of replication quality, underscoring the challenge of replicating structured research output in a HITL system. Unlike automated systems, our approach relies on user insights and interpretations, making the process subjective and context-dependent.
@@
Requirements. ExtracTable satisfies most core requirements defined for the HITL framework. It guides users from defining a research interest to extracting and editing data, with CSV export for seamless integration into scholarly KGs like the ORKG. The modular pipeline supports extensibility, adaptability, and long-term semantic preservation. However, key limitations remain. Entity linking (FR12) is not yet implemented, leaving extracted entities disconnected from existing KG resources. Incorporating entity alignment and exporting in formats like JSON-LD or RDF would improve semantic interoperability. Additionally, performance degrades with large datasets (NR4), as processing occurs sequentially per document. Future work should focus on performance optimization and parallelization strategies. Participants also requested more control over the data model (FR8). Allowing users to define domain-specific properties and extraction instructions would improve precision and adaptability across fields. This aligns with feedback on the need for more customizable and explainable outputs. Crucially, while automation accelerates literature processing, LLM-based extractions still require human validation (FR10, FR11). Without critical oversight, there’s a risk of misinterpreting or blindly trusting generated results. This underscores the ethical implications of AI-assisted research, especially risking the loss of genuine understanding and knowledge internalization.
@@
Future Work. Future development should allow storing and refining outputs with customizable data models, enabling users to iteratively improve extraction accuracy. Users can selectively re-run extractions on specific cells while preserving high-quality results, enhancing the HITL process. Semantic structuring, like linking content to knowledge graph entities and supporting interoperable formats, will strengthen the system’s utility. Ongoing improvements in scalability and model configurability will stabilize and enrich the tool. Lastly, research should focus on user guidance for validating and interpreting extractions, e.g., by highlighting source passages in PDFs, ensuring human oversight in AI-driven workflows.
@@ 7 Conclusion
Our work contributes ExtracTable, a Human-in-the-Loop (HITL) approach for structuring scientific content, demonstrated in the context of the ORKG to showcase its suitability for downstream knowledge graph integration. Aligned with the Quality Improvement (QIP) paradigm, ExtracTable comprises a two-stage workflow — Creating a scientific corpus of literature and Knowledge extraction and validation — implemented as a modular framework in a Jupyter notebook environment. This framework assists researchers in identifying relevant publications, extracting key information, and structuring knowledge into reusable formats ready for integration into scientific knowledge graphs.
By automating repetitive tasks and enabling researchers to interactively refine outputs, ExtracTable reduces the manual effort typically involved in processing large volumes of literature. It moves toward making large-scale literature reviews more efficient and helps researchers keep pace with the rapidly expanding body of scientific knowledge. While evaluation results showed the tool is highly usable and accelerates the process, we identified areas for improvement. Additional support is needed to engage users in the knowledge extraction process, alongside feedback loops to further enhance the quality of extracted information.
Acknowledgements
The authors would like to thank the Federal Government and the Heads of Government of the Länder, as well as the Joint Science Conference (GWK), for their funding and support within the framework of the NFDI4Ing and NFDI4DataScience consortia. This work was partially funded by the German Research Foundation (DFG) - project number 442146713 and project number 460234259.